OpenStudy (anonymous):

Give a medal ? 4* e^x = 17

OpenStudy (jdoe0001):

\(\bf lnx = log_ex\)

OpenStudy (campbell_st):

divide both sides by 4 then take the base e log of both sides

OpenStudy (jdoe0001):

\(\bf 4e^x=17 \implies e^x= \cfrac{17}{4}\\ log_ee^x = log_e\left(\cfrac{17}{4}\right) \implies x = ln\left(\cfrac{17}{4}\right)\\ x = ln(17)-ln(4) \)